# FisheyeDepth: A Real Scale Self-Supervised Depth Estimation Model for Fisheye Camera

### v1

@@ I INTRODUCTION
Depth estimation plays a crucial role in 3D scene perception in the fields of robotics and autonomous driving. Due to the sparse nature and high cost of LiDAR point cloud-based depth sensing, image-based perception methods hold significant value in terms of coverage density and redundancy. Fisheye cameras, with their wide field of view (FOV) and advanced geometric properties, have seen a notable increase in use for perception tasks in recent years [1, 2]. However, the majority of research have focused on pinhole cameras for depth estimation [3, 4, 5, 6], while research dedicated to fisheye cameras remains relatively scarce [7, 8].
Researchers have conducted extensive studies on depth estimation using learning-based methods with narrow FOV cameras [9, 10, 11], but these efforts primarily focus on traditional 2D content captured by cameras, adhering to the typical pinhole projection model based on rectified image sequences. With the rise of affordable wide-angle fisheye cameras, depth estimation has been extended to omnidirectional content through methods like omnidirectional stereo [12, 13] and Structure from Motion (SfM) [14, 15]. However, these approaches often rely on LiDAR or high-precision maps to generate dense, pixel-wise labels for ground truth data, such as in the KITTI dataset [16]. This necessitates multiple LiDAR sensors to compensate for blind spots at close ranges, making the overall setup both expensive and time-consuming, thus limiting the amount of data available for model training. Furthermore, when projecting LiDAR point clouds into pixel-wise distance labels for fisheye cameras, significant motion distortions often occur.
@@ II-A Monocular Self-Supervised Depth Estimation
Several approaches have also developed dedicated self-supervised depth estimation frameworks [29, 30], such as leveraging teacher-student learning models for training [31], with refined depth estimation strategies applied during testing [32, 33]. Moreover, some methods predict camera parameters [24, 34] to facilitate cross-camera training using images captured by different cameras. While all these methods have demonstrated remarkable success on pinhole camera images, they have rarely been adapted to more complex camera models, such as fisheye cameras.
@@ II-B Fisheye Self-Supervised Depth Estimation
Although these methods mitigate some fisheye image distortion through improved models and frameworks, they do not fully address the geometric transformation challenges. Severe distortion, especially at the image edges, still affects depth estimation. Additionally, the lack of real motion scale consideration during training limits their use in real-world robotic applications.
@@ II-C Fisheye Depth Estimation Dataset
The KITTI-360 [39] dataset is one of the few comprehensive fisheye image datasets that is publicly available. It provides rich sensor data and detailed annotations, including dense semantic and instance annotations for both 3D point clouds and 2D images. The dataset was collected using 180° fisheye cameras mounted on both sides of the vehicle, along with a top-mounted push-broom LiDAR scanner, enabling a complete 360° view. Notably, KITTI-360 includes accurate depth annotations, making it a valuable resource for depth estimation tasks.
The WoodScape [40] dataset is designed to advance the application of fisheye cameras in low-speed vehicle scenarios. It features four surround-view cameras and provides instance-level semantic annotations for over 10,000 images, covering nine tasks including segmentation, depth estimation, and 3D bounding box detection. However, WoodScape does not public the ground truth of fisheye image depth, which is a significant limitation for quantitative evaluation.

### v2

@@ I INTRODUCTION
Depth estimation plays a crucial role in 3D scene perception in the fields of robotics and autonomous driving. Due to the sparse nature and high cost of LiDAR point cloud-based depth sensing, image-based perception methods hold significant value in terms of coverage density and redundancy [1, 2]. Fisheye cameras, with their wide field of view (FOV) and advanced geometric properties, have seen a notable increase in use for perception tasks in recent years [3, 4]. However, the majority of research have focused on pinhole cameras for depth estimation [5, 6, 7, 8], while research dedicated to fisheye cameras remains relatively scarce [9].
Researchers have conducted extensive studies on depth estimation using learning-based methods with narrow FOV cameras [10, 11, 12], but these efforts primarily focus on traditional 2D content captured by cameras, adhering to the typical pinhole projection model based on rectified image sequences. With the rise of affordable wide-angle fisheye cameras, depth estimation has been extended to omnidirectional content through methods like omnidirectional stereo [13, 14] and Structure from Motion (SfM) [15, 16]. However, these approaches often rely on LiDAR or high-precision maps to generate dense, pixel-wise labels for ground truth data, such as in the KITTI dataset [17]. This necessitates multiple LiDAR sensors to compensate for blind spots at close ranges, making the overall setup both expensive and time-consuming, thus limiting the data available for model training [18]. Furthermore, when projecting LiDAR point clouds into pixel-wise distance labels for fisheye cameras, significant motion distortions often occur.
@@ II-A Monocular Self-Supervised Depth Estimation
Several approaches have also developed dedicated self-supervised depth estimation frameworks [31, 32], such as leveraging teacher-student learning models for training [33], with refined depth estimation strategies applied during testing [34, 35]. Moreover, some methods predict camera parameters [26, 36] to facilitate cross-camera training using images captured by different cameras. While all these methods have demonstrated remarkable success on pinhole camera images, they have rarely been adapted to more complex camera models, such as fisheye cameras [37].
@@ II-B Fisheye Self-Supervised Depth Estimation
Although these methods mitigate some fisheye image distortion through improved models and frameworks, they do not fully address the geometric transformation challenges [42]. Severe distortion, especially at the image edges, still affects depth estimation. Additionally, the lack of real motion scale consideration during training limits their use in real-world robotic applications.
@@ II-C Fisheye Depth Estimation Dataset
The KITTI-360 [43] dataset is one of the few comprehensive fisheye image datasets that is publicly available. It provides rich sensor data and detailed annotations, including dense depth, semantic, and instance annotations for both 3D point clouds and 2D images. The dataset was collected using 180° fisheye cameras mounted on both sides of the vehicle, along with a top-mounted push-broom LiDAR scanner, enabling a complete 360° view, making it a valuable resource for fisheye depth estimation tasks.
The WoodScape [44] dataset is designed to advance the application of fisheye cameras in low-speed vehicle scenarios. It features four surround-view cameras and provides instance-level semantic annotations for over 10,000 images, covering nine tasks including segmentation, depth estimation, and 3D bounding box detection. However, WoodScape does not public the ground truth of fisheye image depth, which is a limitation for quantitative evaluation.